The exact simple interest of 5000 invested from February 28, 2000 to December 31, 2000 is 1000. What is the rate of interest?

1
Expert's answer
2022-01-19T16:14:39-0500

28.02.2002-31.12.2000 - 307 days


"FV=PV(1+\frac{307}{3663}i)"

PV=5000

FV=5000+1000=6000

"6000=5000(1+\frac{307}{3663}i)"

"1.2=1+\frac{307}{3663}i"

"0.2=\frac{307}{3663}i"


i=0.2384 or 23.84%